# Cluster-based Graph Collaborative Filtering

### v1

@@ 4. Experiments
RQ2: How do the number of graph convolution layers affect ClusterGCF?
@@ 4.1. Experimental setup
We used Tensorflow  [^6] https://www.tensorflow.org. to implement ClusterGCF. For all models, the embedding size was fixed at 64 and initialized using Xavier (Xavier and Yoshua, 2010). For optimization, we used Adam (Kingma and Ba, 2015) with a default learning rate of 0.001. The mini-batch size is 1024 on Office and Kindle Store. In contrast, it is increased to 2048 on Gowalla and Yelp208 to accelerate the model training. The $L_{2}$ regularization coefficient $\lambda$ is searched in $\{1e^{-6},1e^{-5},\cdots,1e^{-1}\}$. The embedding propagation layer is searched in the range of $\{1,2,\cdots,8\}$. The GCMC method uses a single convolutional layer as its original setting. For ClusterGCF and IMP-GCN, the cluster number is searched in the range of $\{2,3,4\}$. Besides, the temperature coefficient is tuned amongst $\{1e^{-2},1e^{-1},\cdots,1e^{2}\}$. The early stopping and validation strategies used are identical to those employed in IMP-GCN. ensuring that we carefully tuned the key parameters.
@@ 4.3. Effect of Graph Convolution Layers (RQ2)
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
In this section, we evaluated the performance of our ClusterGCF model by varying the number of layers in the graph convolution process. Our primary goal is to examine whether the ClusterGCF model can mitigate the over-smoothing problem often encountered in GCN-based recommendation models. To investigate the effectiveness of ClusterGCF, we iteratively increased the number of convolution layers from 1 to 8. We then compared its performance against four GCN-based methods: LightGCN, IMP-GCN, NCL, and GTN. The comparative outcomes are depicted in Figure 4. In particular, both ClusterGCF and IMP-GCN attained their optimal performance while performing first-order graph convolution across the entire graph. IMP-GCN performs high-order graph convolution within the subgraphs (see Section 4.4) while ClusterGCF operates within the cluster-specific graphs. As a result, both two models adopt the same message-passing strategy as LightGCN in the first-order convolution layer. From the results, we had some interesting observations.
First, all three models (LightGCN, NCL, and GTN) achieve their peak performance when stacking 3 or 4 layers across all datasets. Beyond this, they will encounter dramatic performance degradation with increasing layers. These results indicate that these three models suffer from the over-smoothing problem and inadvertently introduce noise information in the deep structure. While both GTN and NCL achieve better performance than LightGCN at all layers. Moreover, they outperform IMP-GCN when stacking less than 3 or 4 layers. This is because GTN reduces unreliable information by capturing the adaptive reliability of the interactions, and NCL benefits from the semantic and structural neighbors among users/items in representation learning.
@@ 4.4. Effect of Cluster-specific Graphs (RQ3)
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
@@ 4.4.2. The Starting Layer of Cluster-based Graph Convolution
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
As illustrated in Figure 6, ClusterGCF exhibits faster convergences than IMP-GCN. For example, ClusterGCF reaches peak performance at 320 epochs, while IMP-GCN requires 360 epochs on the Kindle Store dataset. This can be attributed to the soft clustering method employed in ClusterGCF. Soft clustering provides more flexibility for the model. When the model tries to find the optimal classification or clusters, it does not need to forcefully assign a node to a specific cluster but can instead assign a probability to it. This means that the model can adjust its decision boundaries more smoothly, accelerating the learning process. Due to the smooth nature of soft clustering, the model might encounter fewer local optima and traps. This means that the optimization algorithm could more easily find a good solution. In summary, soft clustering offers smoother and more informative feedback, aiding the model in learning and converging more rapidly.
Refer to captionRefer to caption
Refer to captionRefer to caption

### v2

@@ 4. Experiments
RQ2: How do the number of graph convolution layers impact ClusterGCF?
@@ 4.1. Experimental setup
We used Tensorflow  [^6] https://www.tensorflow.org. to implement ClusterGCF. For all models, the embedding size was fixed at 64 and initialized using Xavier (Xavier and Yoshua, 2010). For optimization, we used Adam (Kingma and Ba, 2015) with a default learning rate of 0.001. The mini-batch size is 1024 on Office and Kindle Store. In contrast, it is increased to 2048 on Gowalla and Yelp208 to accelerate the model training. The $L_{2}$ regularization coefficient $\lambda$ is searched in $\{1e^{-6},1e^{-5},\cdots,1e^{-1}\}$. The embedding propagation layer is searched in the range of $\{1,2,\cdots,8\}$. The GCMC method uses a single convolutional layer as its original setting. For ClusterGCF and IMP-GCN, the cluster number is searched in the range of $\{2,3,4\}$. Besides, the temperature coefficient is tuned amongst $\{1e^{-2},1e^{-1},\cdots,1e^{2}\}$. The early stopping and validation strategies used are identical to those employed in IMP-GCN, ensuring that we carefully tuned the key parameters.
@@ 4.3. Effect of Graph Convolution Layers (RQ2)
In this section, we evaluated the performance of our ClusterGCF model by varying the number of layers in the graph convolution process. Our primary goal is to examine whether the ClusterGCF model can mitigate the over-smoothing problem often encountered in GCN-based recommendation models. To investigate the effectiveness of ClusterGCF, we iteratively increased the number of convolution layers from 1 to 8. We then compared its performance against four GCN-based methods: LightGCN, IMP-GCN, NCL, and GTN. The comparative outcomes are depicted in Fig. 4. In particular, both ClusterGCF and IMP-GCN attained their optimal performance while performing first-order graph convolution across the entire graph. IMP-GCN performs high-order graph convolution within the subgraphs (see Section 4.4) while ClusterGCF operates within the cluster-specific graphs. As a result, both two models adopt the same message-passing strategy as LightGCN in the first-order convolution layer. From the results, we had some interesting observations.
First, all three models (LightGCN, NCL, and GTN) achieve their peak performance when stacking 3 or 4 layers across all datasets. Beyond this, they will encounter dramatic performance degradation with increasing layers. These results indicate that these three models suffer from the over-smoothing problem and inadvertently introduce noise information in the deep structure. Both GTN and NCL achieve better performance than LightGCN at all layers. Moreover, they outperform IMP-GCN when stacking less than 3 or 4 layers. This is because GTN reduces unreliable information by capturing the adaptive reliability of the interactions, and NCL benefits from the semantic and structural neighbors among users/items in representation learning.
@@ 4.4.2. The Starting Layer of Cluster-based Graph Convolution
As illustrated in Figure 6, ClusterGCF exhibits faster convergences than IMP-GCN. For example, ClusterGCF reaches peak performance at 320 epochs, while IMP-GCN requires 360 epochs on the Kindle Store dataset. This can be attributed to the soft clustering method employed in ClusterGCF. Soft clustering provides more flexibility for the model. When the model tries to find the optimal classification or clusters, it does not need to forcefully assign a node to a specific cluster but can instead assign a probability to it. In other words, the model can adjust its decision boundaries more smoothly, accelerating the learning process. Due to the smooth nature of soft clustering, the model might encounter fewer local optima and traps. Hence, the optimization algorithm could more easily find a good solution. In summary, soft clustering offers smoother and more informative feedback, aiding the model in learning and converging more rapidly.
@@ 4.6. Visualization (RQ5)
4.7. Case Study
Figure 9. An example of high-order neighbor nodes in two cluster-specific graphs. The weights associated with each node are marked in red (cluster 1) and blue (cluster 2).
As shown in Fig. 9, we conducted a case study to demonstrate the effectiveness of our proposed method in filtering out noisy neighborhoods. We choose a user node $u_{126}$ at the second graph layer and the interacted three item nodes $i_{10}$, $i_{67}$ and $i_{1093}$ on Office. The weights associated with each node within two cluster-specific graphs are marked in red (cluster 1) and blue (cluster 2). Although the user $u_{126}$ interacted with three items, two of them ($i_{10}$ and $i_{67}$) are assigned greater weights in cluster 1, while the other item ($i_{1093}$) is assigned a greater weight in cluster 2. Consequently, during the embedding propagation process, the influence of this node’s information is reduced in both cluster-specific graphs, demonstrating that our method can effectively filter out noisy information.